OpenStudy (anonymous):

The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work

OpenStudy (anonymous):

I got d=ut + 0.5at^2

OpenStudy (anonymous):

0.5at^2 = d - ut at^2 = 2*(d - ut) t^2 = 2/a*(d-ut) t = \[\sqrt{2/a*(d-ut)}\]

OpenStudy (anonymous):

why is there a square root symbol